010 10.0 points A coin is placed 25 cm from the center of a horizontal turntable, initially at rest. The turntable then begins to rotate. When the speed of the coin is 110 cm/s (rotating at a constant rate), the coin just begins to slip The acceleration of gravity is 980 cm/s What is the coefficient of static friction be-
Answer #1

The coin will begin to slip when the centrifugal force equals friction force

Let mass of coin be m

centrifugal force=m*110²/25=484m

Let coefficient of friction be n

Friction force=n*m*980

484m=n*m*980

n=0.49